# Multi-level Temporal-channel Speaker Retrieval for Zero-shot Voice Conversion

### v2

@@ V-C2 Speaker Retrieval Process
As described in Section III-B, retrieving speaker information in temporal and channel dimensions with multiple granularities is crucial for speaker modeling in zero-shot voice conversion. Therefore, we further investigate the speaker retrieval process in temporal and channel dimensions at multi-level TCR blocks. The attention map is a vital component in the retrieval process. Hence, we first plot the attention map $A_{t_{l}}$ and $A_{c_{l}}$ learned in temporal and channel retrieval from low-level TCR block to high-level TCR block. To be specific, we use one speech utterance as the input of the MTCR speaker module to obtain the attention map. Fig. 9 shows the spectrogram of the speech utterance, corresponding pitch contour, and the attention map. In Fig. 9 (c) and (d), the red line in the horizontal axis marks the boundary of each segment. Note that the attention map of each temporal or channel segment has the dimension of $1\times\gamma_{t}$ or $1\times\gamma_{c}$. For visualization, we flat and repeat the $A_{t_{l}}$ and $A_{c_{l}}$ to the origin temporal and channel length, respectively. The vertical axis value ranges from 0 to 1 and represents the proportion of each feature in the temporal or channel dimensions within a segment. This proportion ensures that the sum of weights within each segment is 1. As can be seen, in the low-level block, the fluctuation of attention weight in both temporal and channel retrieval is relatively small, since the adjacent temporal and channel features within a segment are similar. As the layer deepens, the attention weight value begins to change sharply, indicating that the model is more sensitive to the channel and temporal regions that are highly correlated to the speaker’s SV embedding (e.g. speaker timbre information). For instance, in $A_{t_{2}}$, the attention weight increases at the fifth frame when the speaker begins to speak. In the third segment, from frame 32 to 48, compared with the unvoiced part, the voiced sound (from 44 to 48) with higher attention weight provides more speaker information for the model. And in voice sound from frame 48 to 64, the attention weights change according to the correlation to the speaker timbre. These observations suggest that the speaker retrieval process enables the model to flexibly adjust its attention across temporal and channel regions, considering variations in the amount of speaker information present.
TABLE V: Validation results of the speaker retrieval process.
@@
To further quantify the effect of the speaker retrieval process, we average the attention map in a specific retrieval process, while keeping the other retrieval unchanged. For example, to verify the retrieval process in the first TCR block, only the attention weights of the first block are set to the fixed average value, and the whole modified model is trained from scratch. As shown in Table. V, we implement five variants to verify the efficiency of the retrieval process. Experiments are conducted on 8 unseen speakers mentioned above, and the validation results of the speaker retrieval process are shown in the table. The results show that discarding the retrieval process in any one of the TCR blocks leads to a degradation of performance on WER and ACC. Compared with the retrieval process in block 2 and block 3, averaging the attention map in block 2 shows a relatively obvious degradation in speaker accuracy which indicates the retrieval process in block 2 is important for speaker modeling. Moreover, as shown in the fifth and sixth columns of Table V, averaging attention maps in temporal retrieval or channel retrieval in all blocks apparently shows an increase in WER and a decrease in speaker accuracy. This result suggests that temporal and channel retrieval are helpful for capturing speaker timbre. Interestingly, changing temporal retrieval has a relatively large degradation in intelligibility, because the attention map $A_{t_{l}}$ learned in the temporal retrieval is used for content-based alignment in the fusion block, as shown in Fig. 2.
@@ VI Conclusion
In this paper, we introduce MTCR-VC, a new zero-shot VC model with a novel speaker modeling approach. Specifically, we propose the MTCR speaker module, which hierarchically queries speaker timbre from the target speaker’s speech across different temporal and channel granularities by stacking multiple TCR blocks. This allows the model for flexible allocation of attention across temporal and channel regions, based on the richness of speaker information at different stages, thus improving the robustness of speaker timbre modeling. Moreover, applying the cycle-based training strategy facilitates the process of speech reconstruction and disentanglement toward the aim of the zero-shot VC task. Our experimental results show that the proposed MTCR-VC achieves superior performance in modeling speaker timbre while maintaining good speech naturalness.
Although the experimental results have demonstrated good performance of the proposed method on speaker timbre modeling, we have to point out that there are still some limitations. Specifically, the zero-shot VC performance, including naturalness and speaker similarity, varies among different speakers and speeches with different recording conditions, likely due to the limited speaker diversity and recording environments in the training corpus. To address this limitation, future work should consider modeling speaker timbre from more diverse speech data, containing even a larger number of speakers and speech recordings, to improve the performance and generalization of zero-shot VC.

### v3

@@ V-C2 Speaker Retrieval Process
As described in Section III-B, retrieving speaker information in temporal and channel dimensions with multiple granularities is crucial for speaker modeling in zero-shot voice conversion. Therefore, we further investigate the speaker retrieval process in temporal and channel dimensions at multi-level TCR blocks. The attention map is a vital component in the retrieval process. Hence, we first plot the attention map $A_{t_{l}}$ and $A_{c_{l}}$ learned in temporal and channel retrieval from low-level TCR block to high-level TCR block. Specifically, we use one speech utterance as the input of the MTCR speaker module to obtain the attention map. Fig. 9 shows the spectrogram of the speech utterance, corresponding pitch contour, and the attention map. In Fig. 9 (c) and (d), the red line in the horizontal axis marks the boundary of each segment. Note that the attention map of each temporal or channel segment has the dimension of $1\times\gamma_{t}$ or $1\times\gamma_{c}$. For visualization, we flat and repeat the $A_{t_{l}}$ and $A_{c_{l}}$ to the origin temporal and channel length, respectively. The vertical axis value ranges from 0 to 1 and represents the proportion of each feature in the temporal or channel dimensions within a segment. This proportion ensures that the sum of weights within each segment is 1. As can be seen, in the low-level block, the fluctuation of attention weight in both temporal and channel retrieval is relatively small since the adjacent temporal and channel features within a segment are similar. As the layer deepens, the attention weight value begins to change sharply, indicating that the model is more sensitive to the channel and temporal regions that are highly correlated to the speaker’s SV embedding (e.g. speaker timbre information). For instance, in $A_{t_{2}}$, the attention weight increases when the speaker begins to speak in the fifth frame. In the third segment, from frame 32 to 48, compared with the unvoiced part, the voiced sound (from 44 to 48) with higher attention weight provides more speaker information for the model. In voice sound from frames 48 to 64, the attention weights change according to the correlation to the speaker timbre. These observations suggest that the speaker retrieval process enables the model to flexibly adjust its attention across temporal and channel regions, considering variations in the amount of speaker information present.
TABLE VI: Validation results of the speaker retrieval process.
@@
To further quantify the effect of the speaker retrieval process, we average the attention map in a specific retrieval process while keeping the other retrieval unchanged. For example, to verify the retrieval process in the first TCR block, only the attention weights of the first block are set to the fixed average value, and the whole modified model is trained from scratch. As shown in Table. VI, we implement five variants to verify the efficiency of the retrieval process. Experiments are conducted on 8 unseen speakers mentioned above, and the validation results of the speaker retrieval process are shown in the table. The results show that discarding the retrieval process in any one of the TCR blocks leads to a degradation of performance on WER and ACC. Compared with the retrieval process in block 2 and block 3, averaging the attention map in block 2 shows a relatively obvious degradation in speaker accuracy, which indicates that the retrieval process in block 2 is important for speaker modeling. Moreover, as shown in the fifth and sixth columns of Table VI, averaging attention maps in temporal retrieval or channel retrieval in all blocks apparently shows an increase in WER and a decrease in speaker accuracy. This result suggests that temporal and channel retrieval help capture speaker timbre. Interestingly, changing temporal retrieval has a relatively large degradation in intelligibility because the attention map $A_{t_{l}}$ learned in the temporal retrieval is used for content-based alignment in the fusion block, as shown in Fig. 2.
@@ VI Conclusion
This paper introduces MTCR-VC, a new zero-shot VC model with a novel speaker modeling approach. Specifically, we propose the MTCR speaker module, which hierarchically queries speaker timbre from the target speaker’s speech across different temporal and channel granularities by stacking multiple TCR blocks. This allows the model for flexible allocation of attention across temporal and channel regions based on the richness of speaker information at different stages, thus improving the robustness of speaker timbre modeling. Moreover, applying the cycle-based training strategy facilitates the process of speech reconstruction and disentanglement toward the aim of the zero-shot VC task. Our experimental results show that the proposed MTCR-VC performs better in modeling speaker timbre while maintaining good speech naturalness.
Although the experimental results have demonstrated good performance on speaker timbre modeling, we have to point out that there are still some limitations. Specifically, the zero-shot VC performance, including naturalness and speaker similarity, varies among different speakers and speech recordings with different recording conditions, likely due to the limited speaker diversity and recording environments in the training corpus. To address this limitation, future work should consider modeling speaker timbre from more diverse speech data, containing even a larger number of speakers and speech recordings, to improve the generalization ability of zero-shot VC.